# Static or Dynamic: Towards Query-Adaptive Token Selection for Video Question Answering

### v1

@@ 5 Experiments
5.1 Experiment Settings
Model Settings EgoSchema VideoMME
@@ 5.1 Experiment Settings
Qwen2-VL-7B Original 64 - 66.2 71.1 59.4 50.8 60.4
Retrieval 256 64 63.6 71.0 61.3 52.2 61.5
Similarity 256 64 66.6 71.4 60.6 51.8 61.3
Ours 256 64 67.8 72.4 63.1 53.2 62.9
Original 32 - 64.7 68.9 55.2 48.7 57.6
Retrieval 128 32 61.7 70.0 58.6 51.6 60.0
Similarity 128 32 65.6 70.1 58.7 51.8 60.2
Ours 128 32 66.7 71.4 61.0 51.7 61.4
LLaVA-OneVision-7B Original 64 - 60.1 70.6 55.8 47.8 58.0
Retrieval 256 64 57.7 64.0 53.4 47.0 54.8
Similarity 256 64 59.6 71.0 57.9 50.8 59.9
Ours 256 64 60.3 71.9 58.3 51.4 60.6
Original 32 - 60.4 71.3 57.4 48.0 58.9
Retrieval 128 32 57.9 63.2 53.9 46.0 54.4
Similarity 128 32 60.2 70.8 57.1 49.7 59.2
Ours 128 32 60.5 70.2 58.0 51.6 59.9
Table 2: Long video benchmarks results.
  • Benchmarks. To comprehensively evaluate performance, we select benchmarks for both long and short videos. For long videos, we utilize VideoMME and EgoSchema. And for short videos, we employ MSVD-QA.
    • VideoMME Fu et al. (2024a) comprises 900 videos and 2,700 multiple-choice questions across six domains. Videos range from 11 seconds to 1 hour in duration and are categorized as short, medium, and long subsets.
    • EgoSchema Mangalam et al. (2023) features over 5,000 multiple-choice questions. The videos have an average length of about three minutes and cover a wide range of human activities and behaviors. We use its full set.
    • MSVD-QA Xu et al. (2017) contains 1,970 video clips and about 50,500 question-answer pairs. The average video duration is 10 seconds. We used the test set which includes around 13,000 questions.
    We use multiple-choice accuracy as the metric for VideoMME and EgoSchema. For MSVD-QA, which consists of open-ended questions, we use GPT OpenAI (2024) for evaluation and scoring.
  • Baselines. We validate our method on the currently well recognized Qwen2-VL Wang et al. (2024) and LLaVA-OneVision Li et al. (2024a). Qwen2-VL introduces dynamic resolution support and multimodal rotary position embeddings. LLaVA-OneVision is capable of handling single-image, multi-image, and video tasks. Here we adopt the 7B version for both models.
    As shown in Table 1, previous methods either compress within the KV cache, failing to address the issue of long token sequences input to the large language model, or introduce new modules that require training the model, making it unfair to compare them directly with our training-free approach. Thus we select three baselines: 1) Original: uniform sampling frames within the budget; 2) Retrieval: higher-frequency uniform sampling frames beyond the budget, followed by pruning based on cosine similarity with query embeddings to meet the token limit; 3) Similarity: higher-frequency uniform sampling frames beyond the budget, followed by pruning based on the cosine similarity of adjacent token embeddings to meet the token limit. In practice, the idea of “Retrieval” and “Similarity” are often utilized in compression modules Qian et al. (2025); Song et al. (2024); He et al. (2024a).
    To further demonstrate the advantage of our method, we compare it with 4) LongVU Shen et al. (2024), a training-based approach. For fairness, we reproduce its compression method on both models and evaluate it in a training-free setting.
  • Implementation Details. All experiments are conducted on two 40GB A100 GPUs. For multiple-choice questions, the model generates one token, while for open-ended questions, outputs are limited to 30 tokens. The prompts used for each dataset are detailed in Appendix A.
    Note that video resolution affects the number of frame tokens generated by Qwen2-VL. A fixed token budget may result in inconsistent numbers of frames that can be processed across videos, complicating experiments and comparisons. To address this, we convert the token limit into a frame-based constraint, setting the budget frame number as $T_{b}$. For a video generating $H\times W$ tokens per frame, the token budget becomes $L_{b}=T_{b}\times H\times W$. This adjustment simplifies implementation and ensures fair sample comparisons.
5.2 Main Results
  • Long Video Results. Table 2 presents the results on EgoSchema and VideoMME under two settings: 256-frame sampling with 64-frame budget (256-64) and 128-frame sampling with 32-frame budget (128-32). Our method outperforms baselines across both benchmarks and most subsets.
    Qwen2-VL-7B shows significant improvements under the 256-64 setting, surpassing baselines by up to 4.2% on EgoSchema and 2.5% on VideoMME. In the 128-32 setting, it outperforms baselines by up to 5.0% and 3.8%, and achieves a 5.8% improvement over the “Original” method on VideoMME’s medium subset. LLaVa-OneVision-7B achieves 60.3% and 60.6% accuracy on the two benchmarks under the 256-64 setting, and 60.5% and 59.9% under 128-32, showing weaker advantages. This is likely due to noise introduced by sparse token selection under one-dimensional positional encoding. Qwen2-VL-7B, with its three-dimensional positional encoding, exhibits lower noise, highlighting the benefits of specialized positional embedding design for VideoLMs.
    Overall, these results demonstrate the effectiveness of our method, and reveal some model-specific behaviors and limitations.
  • Short Video Results. Table 3 presents the results on MSVD-QA. Given that average video length is only 10 seconds, we reduce the frame sampling rate to 64 frames and adjust the budget to 16 frames. Our method still achieves good performance on Qwen2-VL-7B, surpassing all baselines. It outperforms the “Retrieval” approach by 2.4% in accuracy and nearly 0.1 in score. However, a contrasting trend is observed on LLaVa-OneVision-7B. While our method exceeds both the “Original” and “Similarity” approach, it is slightly outperformed by the “Retrieval” method, with marginal differences. These results highlight that the baselines exhibit instability across different models and datasets, while our approach delivers robust performance, demonstrating its reliability and effectiveness across diverse experimental settings.
    Model Method MSVD-QA
       Acc Score
    Qwen2-VL Original 66.0 3.59
    Retrieval 64.4 3.52
    Similarity 66.5 3.60
    Ours 66.8 3.61
    LLaVA-OneVision Original 54.3 3.09
    Retrieval 54.8 3.12
    Similarity 54.3 3.10
    Ours 54.7 3.11
    Table 3: Short video benchmark results with 64-frame sampling and 16-frame budget.
    Figure 4: In our qualitative analysis cases, our method allocates static and dynamic tokens at a ratio of 3:1 for Question 1 which is a OCR problem, and a ratio of 5:11 for Question 2 which pertains to action recognition.
@@ 5.2 Main Results
    Ours 67.8 62.9
    LLaVA- OneVision Original 60.1 58.0
@@
    Ours 60.3 60.6
    Table 4: Comparison with LongVU under a training-free setting with 256-frame sampling and 64-frame budget.
    Method EgoSchema VideoMME
    Orignal 64.7 57.6
    Explore + Random 66.3 60.7
    Explore + Select 66.7 61.4
    Table 5: Ablation about exploration and selection stage.
    Figure 5: Search space size analysis. The x-axis represents the search space size. There are $n$ subsequences in the space, and thier static frame number ranges from $\{1,2,\dots,n\}$. Assuming the budget frame is $N_{b}$, “$1$” refers to $n=N_{b}$, “$1/2$” indicates $n=\lfloor N_{b}/2\rfloor$, “$1/4$” represents $n=\lfloor N_{b}/4\rfloor$.
  • Comparison with training-free LongVU. Many visual token compression methods require training, such as LongVU Shen et al. (2024). For a fair comparison, we implement its compression component in a training-free manner. Following the original paper, we use DINOv2 Oquab et al. (2023) with a 0.83 threshold for frame reduction, downsampling at a $\lfloor 2/3\rfloor$ ratio. In the final step, we find it difficult to remove tokens to meet budget using the threshold. Therefore, we opt for top-K selection for precise token control. As shown in Table 4, our method outperforms LongVU across models and datasets. Moreover, compared to LongVU, our method offers better control over the token count.
5.3 Ablation Study
As shown in Table 5, we conduct a two-stage ablation study on our method. The ablation experiments were performed on Qwen2-VL-7B, sampling 128 frames with a budget of 32 frames. First, we validated the effectiveness of the exploration stage. As indicated by the “Explore” row in the table, generating multiple token subsequences followed by random selection results in improvement compared to the original operation, demonstrating the rationality of our search space design. Then we verify the effectiveness of the selection phase. On both benchmarks, our selection method achieves improvement over the random selection.
5.4 Further Analysis
  • Qualitative Analysis. Figure 4 shows two questions for the same video, which are correctly answered employing our method. Question 1, an OCR problem predominantly reliant on static information, prompts the method to allocate a static-to-dynamic token ratio of 3:1. Conversely, action-ralated Question 2, necessitating the identification of a player scoring a goal, leads to the adoption of a static-to-dynamic token ratio of 5:11.
  • Search Space Size. We study search space size impact using Qwen2-VL-7B, sampling 128 frames with a 32-frame budget. Figure 5 shows performance improves on EgoSchema when search space matches the budget but declines on VideoMME. We attribute this to excessive static frames, causing sparse dynamic delta token selection and deviation from the training distribution, reducing effectiveness. Additionaly, time cost rises with search space size. To balance these, we set the search space to half the budget frame number.
  • Time Overhead. The compression inevitably incurs time overhead, but our focus is on memory efficiency and information retention. Our overhead mainly comes from metric computation. Using shallow attention layers with question-length queries, sequential processing of 16 candidates in a 28-layer model doubles first-token latency but leaves subsequent decoding unaffected, benefiting open-ended questions. Parallel processing can further reduce latency. On LLaVa-OneVision-7B (128-frame sampling with a 32-frame budget), compared to the 2.24s overhead incurred by LongVU, our approach costs only 0.43s without flash-attention.
6 Conclusion
Given that long videos possess tokens far exceeding the capacity that models can process, we advance token compression strategies by unveiling the following crucial fact: different question types exhibit varying dependencies on dynamic and static information. Based on this discovery, we propose a novel token selection strategy for video token compression. Our method splits video frames into static and dynamic categories, and adaptively determines the optimal token allocations among static and dynamic frames guided by each specific query. Experiments demonstrate the effectiveness and generalizability of our method across multiple models and datasets.
Limitations
In this paper, we propose a novel token selection strategy for token compression in video question answering tasks, addressing the varying dependencies of questions on dynamic and static video information. While the effectiveness of our method has been validated across multiple datasets, certain limitations remain. Firstly, due to differences in positional encoding mechanisms across models, some encoding schemes may impact the model’s ability to accurately judge video length and temporally localize events. Nevertheless, we believe our approach holds insight for developing compression modules in pre-trained and fine-tuned video models. Additionally, although our method incurs no additional memory overhead (superior to pruning in the key-value cache), it does introduce time overhead. This overhead mainly stems from metric computation. We utilize the output of a shallow (second-layer) attention mechanism to compute the metric, where only the attention map between query tokens and vision tokens is computed. This overhead only happens during the initial token inference and does not affect subsequent token generation. It is worth noting that such additional time cost is a common challenge for most compression methods.
Appendix A Prompt Details

### v2

@@ 5 Experiments
5.1 Experimental Settings
  • Benchmarks. To comprehensively evaluate performance, we select benchmarks for both long and short videos. We use VideoMME, EgoSchema, and MLVU for long videos, and MSVD-QA and ActivityNet-QA for short videos.
    VideoMME Fu et al. (2024a) contains 900 videos (11 seconds to 1 hour) and 2,700 QA pairs. EgoSchema Mangalam et al. (2023) includes over 5,000 questions based on videos averaging 3 minutes in length. MLVU Zhou et al. (2024) provides videos ranging from 3 minutes to 2 hours, with the test set containing over 500 QA pairs. MSVD-QA Xu et al. (2017) includes 1,970 short clips (10 seconds on average), with a test split of approximately 13,000 questions. ActivityNet-QA Yu et al. (2019) provides 800 videos and 8,000 QA pairs in the test set.
    We adopt multiple-choice accuracy as the evaluation metric for VideoMME, EgoSchema, and MLVU, and employ GPT-4o mini OpenAI (2024) to score answers for the open-ended MSVD-QA and ActivityNet-QA.
  • Baselines. We validate our plug-and-play method on two representative models: Qwen2-VL Wang et al. (2024), featuring dynamic resolution and multimodal rotary position embeddings, and LLaVA-OneVision Li et al. (2024a), supporting multiple tasks, both in their 7B versions. Results for Qwen2.5-VL Bai et al. (2025) are included in Appendix A.1.
    As shown in Table 1, prior methods either compress only within the KV cache, leaving long input sequences unaddressed, or require training models, making direct comparison with our training-free approach unfair. Thus, we consider three baselines: 1) Original: uniform frame sampling within the token budget; 2) Retrieval: oversample frames, then prune based on cosine similarity between frame and query embeddings to fit the token limit; 3) Similarity: oversample frames, then prune based on cosine similarity between adjacent token embeddings. In practice, both “Retrieval” and “Similarity” strategies are commonly adopted in compression modules Qian et al. (2025); Song et al. (2024); He et al. (2024a). For reference, we also report results from several training-based video understanding methods Li et al. (2023b); Zhang et al. (2024b); Ye et al. (2024); Shen et al. (2024) in the first block of Table 2, though they are not directly comparable due to training cost differences. To further validate the advantages of our method, we include a comparison with our reproduced training-free LongVU in Appendix A.2.
  • Implementation Details. All experiments are run on two 40GB A100 GPUs or one 80GB A100 GPU. For multiple-choice questions, the model generates one token (three for MLVU), while for open-ended questions, outputs are limited to 30 tokens. The prompts used are detailed in Appendix B. Sampling is disabled to ensure deterministic results.
    Note that video resolution affects the number of frame tokens generated by Qwen2-VL, making a fixed token budget yield varying frame counts across videos and complicating comparisons. To address this, we set a frame-based budget $T_{b}$, so the token limit is $L_{b}=T_{b}\times H\times W$, where $H\times W$ is the token count per frame. This approach streamlines implementation and ensures fair comparison. Besides, unless otherwise specified, the number of subsequences generated during the exploration stage is set to half of the frame budget $T_{b}$.
5.2 Main Results
  • Long Video Results. Table 2 shows results on long video benchmarks for two settings: 256-frame sampling with a 64-frame budget (256-64) and 128-frame sampling with a 32-frame budget (128-32). Our method outperforms baselines across all benchmarks and most subsets. Qwen2-VL-7B significantly outperforms baselines by up to 4.2% on EgoSchema, 2.5% on VideoMME, and 5.0% on MLVU (256-64), and by up to 5.0%, 3.8%, and 5.4% (128-32), with a 5.8% gain on VideoMME medium subset. While our method also achieves notable improvements on LLaVA-OneVision-7B, the gains are less pronounced than on Qwen2-VL, likely due to noise from its one-dimensional positional encoding. The three-dimensional positional encoding of Qwen2-VL-7B offers more stable results, highlighting the importance of positional encoding design. Overall, these results demonstrate the effectiveness of our method and reveal some model-specific behaviors and limitations.
  • Short Video Results. Short video benchmarks inherently contain fewer frames, simpler scenes, and primarily coherent motion, making them less affected by token length limitations. As a result, the trade-off between static and dynamic information is less pronounced, and performance gains tend to be smaller compared to long video settings. Nonetheless, we evaluate our method’s generalization on short video benchmarks by sampling 64 frames and setting the budget to 16 for videos averaging 10 seconds. As shown in Table 3, our method consistently outperforms all baselines on Qwen2-VL-7B, achieving up to 3.8% higher accuracy and 0.16 higher scores. On LLaVA-OneVision-7B, it achieves strong results on ActivityNet-QA and performs comparably to the “Retrieval” baseline on MSVD-QA. These results demonstrate the robustness and generalization ability of our method even under short video scenarios.
    Model Method MSVD-QA ActivityNet-QA
    Acc Score Acc Score
    Qwen2-VL Original 66.0 3.59 50.3 2.82
    Retrieval 64.4 3.52 48.6 2.74
    Similarity 66.5 3.60 51.4 2.87
[…]
    LLaVA-
    OneVision
    Original 54.3 3.09 52.6 2.90
    Retrieval 54.8 3.12 50.1 2.77
    Similarity 54.3 3.10 52.4 2.89
[…]
    Table 3: Results on short video benchmarks. Although primarily focused on long videos, our method show stable and generalizable performance on short videos.
    Figure 4: The qualitative analysis demonstrates that our method adjusts token allocation according to the query. For Question 1, an OCR task, the ratio of key-frame tokens to delta-frame tokens was 3:1. In contrast, for Question 2, an action recognition task, the ratio was 5:11.
    Method EgoSchema VideoMME MLVU
    Original 64.7 57.6 46.8
    Explore + Random 66.3 60.7 50.2
    Explore + Select 66.7 61.4 52.2
    Table 4: Ablation study of our method. Results demonstrate the effectiveness of both stages, with each component yielding improvements over the baseline.
    Model Method EgoSchema VideoMME
    Qwen2-VL
    w/ query 66.3 61.6
    w/o query 66.7 61.4
    mean 66.0 60.9
    max 66.7 61.4
    Table 5: Ablation study on metric design. The first block shows that including the query token in $\bm{K}$ has a negligible impact, so it is omitted. The second block finds that the max operation in Equation (9) outperforms the mean on both benchmarks.
    Figure 5: Search space size analysis. The x-axis represents the search space size. There are $n$ subsequences in the space, and their key frame number ranges from $\{1,2,\dots,n\}$. Assuming the budget frame is $N_{b}$, “11” refers to $n=N_{b}$, “1/21/2” indicates $n=\lfloor N_{b}/2\rfloor$, “1/41/4” represents $n=\lfloor N_{b}/4\rfloor$. Larger search spaces benefit EgoSchema but hurt VideoMME and increase the time cost. A balanced setting uses half the budget size.
5.3 Ablation Studies
  • Stage Ablation. As shown in Table 4, we conduct a two-stage ablation study on our method. The ablation experiments are performed on Qwen2-VL-7B, sampling 128 frames with a budget of 32 frames. First, we validate the effectiveness of the exploration stage. As indicated by the “Explore + Random” row in the table, generating multiple token subsequences followed by random selection results in improvement compared to the original operation, demonstrating the rationality of our search space design. Then we verify the effectiveness of the selection phase. On all benchmarks, our selection method achieves improvement over random selection.
  • Metric Ablation. Table 5 presents two ablation studies on our metric design using Qwen2-VL-7B (128-frame sampling, 32-frame budget). The first block compares including or excluding the query token in the construction of $\bm{K}$ in Equation (7), finding only marginal differences; for simplicity, we exclude the query token in our final design. The second block compares max and mean operations for query aggregation in Equation (9), showing that the max operation consistently yields better results, thus supporting our metric choice.
5.4 Further Analysis
  • Qualitative Analysis. Figure 4 shows two questions from the same video, both correctly answered using our method. In this example, we employ 128-frame sampling with a 32-frame budget, and the search space is defined as the full frame budget. Question 1, an OCR problem predominantly reliant on static information, prompts the method to allocate a key-to-delta-frame token ratio of 3:1. Conversely, action-related Question 2, necessitating the identification of a player scoring a goal, leads to the adoption of a key-to-delta-frame token ratio of 5:11.
  • Search Space Size. We investigate the impact of search space size on performance using Qwen2-VL-7B, sampling 128 frames with a 32-frame budget. Figure 5 shows that performance improves on EgoSchema when the search space matches the budget, but declines on VideoMME. We attribute this to excessive key frames, causing sparse delta-frame token selection and deviation from the training distribution, reducing effectiveness. Additionally, the time cost rises with the search space size. Therefore, we set the search space to half the frame budget to balance these factors.
  • Efficiency Trade-off. With a limited token budget, sampling more frames prior to compression and input can improve performance. As shown in Table 6, compressing 32 to 256 sampled frames into a 32-frame budget consistently improves accuracy. While compression introduces computational overhead, our primary focus is on memory efficiency and information retention. This approach involves a trade-off between performance and time cost; all reported time measurements are averaged over 10 runs to reduce variance. The main overhead stems from token selection and the increased load on the vision encoder. The selection cost, introduced by our framework, remains stable at approximately 0.5–0.6s and affects only the first-token latency, leaving subsequent decoding unaffected. The encoding cost scales roughly linearly with the number of sampled frames, but this is a general issue that can be largely mitigated by using asynchronous computation in multi-GPU environments.
    Metric 32 64 128 256 512
    Accuracy (%) 46.8 52.0 52.2 54.0 54.6
    Encoding (s) 0.125 0.222 0.424 0.827 1.628
    Selection (s) - 0.537 0.557 0.565 0.575
    Table 6: Performance and efficiency trade-off on the MLVU dataset using Qwen2-VL. Sampling more frames before compression improves accuracy. While the encoding cost increases roughly linearly (not specific to our method), the selection cost introduced by our framework remains stable.
6 Conclusion
Given that long videos possess tokens far exceeding the capacity that models can process, we advance token compression strategies by unveiling the following crucial fact: different question types exhibit varying dependencies on dynamic and static information. Based on this discovery, we propose a novel token selection strategy for visual token compression. Our method splits video frames into key and delta frames, and adaptively determines the optimal token allocations among key and delta frames guided by each specific query. Experiments demonstrate the effectiveness and generalizability of our method across multiple models and datasets.
Limitations
In this paper, we propose a novel token selection strategy for visual token compression in video question answering, addressing the varying dependencies of questions on dynamic versus static video information. Our method has demonstrated effectiveness across multiple datasets, yet certain limitations remain. First, variations in positional encoding mechanisms across models may affect the ability to accurately estimate video length and temporally localize events. Nonetheless, we believe our approach offers valuable insights for designing compression modules in both pre-trained and fine-tuned video models. In terms of efficiency, our method introduces no additional memory overhead (superior to pruning in the key-value cache) but does incur extra time cost. This overhead mainly arises from the token selection process, where the selection metric is computed using the output of a shallow (second-layer) attention layer. Here, only the attention map between query and visual tokens is calculated, and multiple subsequences must be compared to finalize the selection. Importantly, this overhead occurs only during the initial token inference and does not affect subsequent decoding, and such costs are a common trade-off in most compression methods.
Acknowledgments
This research is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (#023618-00001, RG99/23), and the Start-Up Grant (#023284-00001) of Nanyang Technological University, Singapore.
Appendix A Additional Experiments
The appendix presents additional experiments, including results on the advanced Qwen2.5-VL model and comparisons between our method and the reproduced training-free LongVU approach.
A.1 Experiments on Qwen2.5-VL
Qwen2.5-VL is the latest vision language model in the Qwen series models, officially released in February 2025. Building upon the foundation of Qwen2-VL, Qwen2.5-VL introduces significant enhancements in long video comprehension. Notably, it incorporates absolute time encoding, enabling the model to handle videos of extended durations with second-level event localization. To provide a more comprehensive evaluation of our method, we report experimental results on the Qwen2.5-VL-7B model using the same experimental settings as in the main text.
Model Settings EgoSchema VideoMME MLVU
@@ Appendix A Additional Experiments
Qwen2.5-VL-7B Original 256 - 60.3 75.0 61.8 51.0 62.6 50.0
Retrieval 256 64 60.9 75.4 66.7 54.8 65.6 56.2
Similarity 256 64 60.8 74.0 64.7 54.3 64.3 53.6
[…]
Original 128 - 59.1 73.1 60.0 49.6 60.9 47.2
Retrieval 128 32 60.2 74.6 64.8 53.3 64.2 48.4
Similarity 128 32 60.0 73.3 60.9 51.6 61.9 47.6
[…]
Table 7: Long video benchmark results on Qwen2.5-VL-7B. Our method achieves the best performance on EgoSchema, MLVU, and the VideoMME long subset, with improvements of up to 5.1% on the VideoMME long subset and 8.4% on MLVU over the baselines, demonstrating strong effectiveness and generalization.
Model Method ActivityNet-QA
  Accuracy   Score
Qwen2.5-VL Original 52.1 2.96
Retrieval 52.7 2.98
Similarity 53.1 2.99
[…]
Table 8: Short video benchmark results on Qwen2.5-VL. Our method achieves the highest accuracy and score, outperforming all baselines.
@@
[…]
LLaVA-
OneVision
Original 60.1 58.0
@@
[…]
Qwen2.5-VL Original 60.3 62.6
LongVU 61.6 64.5
[…]
Table 9: Comparison with the reproduced training-free LongVU. Our method consistently outperforms the reproduced LongVU across models and benchmarks, while offering more precise control over the token count.
  • Long Video Results. Table 7 presents the long video benchmark results on Qwen2.5-VL-7B under different sampling and budget settings. Across both the 256-64 and 128-32 configurations, our method consistently achieves the best performance on EgoSchema, MLVU, and the VideoMME long subset. Notably, it improves performance on the VideoMME long subset and MLVU by up to 5.1% and 8.4% compared to the baselines. For the VideoMME short and medium subsets, due to the shorter video length, our approach does not significantly outperform all baselines but still delivers strong results. These findings demonstrate the effectiveness and robustness of our method and further validate its strong generalization capability across different models.
  • Short Video Results. Although our method is primarily designed for long video understanding, it also delivers strong results on short video tasks. For example, on Qwen2.5-VL evaluated with ActivityNet-QA under the 64-frame sampling and 16-frame budget setting, our approach achieves the best performance among all baselines. As shown in Table 8, it attains the highest accuracy of 54.3% and a score of 3.07, outperforming the baselines by up to 2.2% in accuracy and 0.11 in score. These results further demonstrate that our method remains robust and effective across different video lengths.
A.2 Comparison with Training-free LongVU
To further highlight the advantages of our approach, we compare it with LongVU Shen et al. (2024) by reproducing its compression method in a training-free setting under the 256-frame sampling and 64-frame budget configuration. Following the original paper, we use DINOv2 Oquab et al. (2023) with a 0.83 threshold for frame reduction and apply a $\lfloor 2/3\rfloor$ downsampling ratio. However, we observe that achieving an exact token budget with LongVU requires careful tuning of thresholds and heuristics, providing only indirect control over compression. In contrast, our method employs top-K selection, allowing direct and precise control of the token count. As shown in Table 9, our approach consistently outperforms the reproduced LongVU across all models and benchmarks, while offering more reliable and practical token budget management.
Appendix B Prompt Details